# Optimizing Large Language Model Training Using FP4 Quantization

### v1

@@ 1 Introduction
We conduct comprehensive experiments to demonstrate that our FP4 training framework achieves accuracy comparable to models trained in BF16 or FP8 formats with the same hyperparameters. Leveraging the FP8 tensor cores of NVIDIA H100 GPUs to emulate FP4 computations, we train LLMs with up to 13B parameters and 100B training tokens, with minor training loss gap. For zero-shot evaluation on downstream tasks, model trained with FP4 show competitive results against BF16 models. We anticipate better speed performance gains with the availability of next-generation hardware like NVIDIA’s B-series GPUs. We will open-source our training code to facilitate future research and adoption.
@@ 3.1 Differentiable Gradient Estimator
Suppose we quantify the model weight $W$ with a non-differentiable quantization function $f$: $W_{q}=f(W)$. Considering the backward gradient computation for a linear function with quantized weight, the forward pass can be expressed as:
@@
$f(x)=\delta\cdot\big{(}1+\text{sign}(x-\frac{\delta}{2})\cdot|x-\frac{\delta}{2}|^{\frac{1}{k}}\big{)}$ (7)
Figure 3(a) illustrates this function under $k=5$ for the range $[0,0.5]$, which represents the first positive quantization interval in the E2M1 quantization scheme. This figure also shows that under the assumption of STE, forward quantization function is equivalent to $f(x)=x$ because $f^{\prime}(x)\equiv 1$. In Equation 7, $\delta$ represents the quantization interval, and $k$ is a parameter that controls the degree of approximation. As $k$ increases, the function curve becomes sharper and more closely resembles the behavior of the original hard quantization function.
@@
$f^{\prime}(x)=\frac{1}{k}\cdot|x-\frac{\delta}{2}|^{\frac{1}{k}-1}$ (8)
Figure 3(b) and Figure 3(c) show the complete quantization curve $f(x)$ and its derivative $f^{\prime}(x)$ under $k=5$ within the full E2M1 quantization framework. This framework consists of 14 distinct quantization intervals. To prevent excessively large gradient values, the magnitude of $f^{\prime}(x)$ is capped at 3.0, impacting only a very small subset of elements.
In practical model training, the Differentiable Gradient Estimator (DGE) is seamlessly integrated into the process. During the forward pass, we retain the hard quantization function for computational efficiency. For the backward pass, a correction term derived from Equation 8 is applied to the weight gradient calculation following Equation 6. Supplementary integration process and proof for the DGE method in actual training process is provided in Appendix B.
@@ 3.2 Outlier Clamping and Compensation
To address this issue, we propose the Outlier Clamping and Compensation method (OCC) to restrict the range of activation tensors and mitigate the underflow problem. Specifically, we identify outliers—values with the largest absolute magnitudes—through quantile searching and clamp them to a predefined threshold. Given a pre-defined quantile $\alpha$, the clamping function can be expressed as:
@@ 4.1 Experiment Setup
During LLM training, General Matrix Multiplication (GeMM) accounts for over 95% of the computational workload, with this proportion increasing for larger models. Consistent with prior works (Xi et al., 2023; Yang et al., 2020; Dettmers et al., 2022), we focus on 4-bit quantization for GeMM operations, a core feature of FP4 tensor cores. In GeMM $Y=AW$, where $A$ (sequence length × input channels) is the activation tensor and $W$ (input channels × output channels) is the weight tensor, quantization is applied along distinct dimensions to align with matrix multiplication logic: $A$ is quantized token-wise (sequence length dimension), while $W$ is quantized channel-wise (output channels dimension). The aforementioned accuracy-preserving techniques are integrated to minimize quantization error. Since FP4 Tensor Cores are unavailable, we validate FP4 performance using Nvidia H-series GPUs’ FP8 Tensor Cores, which encompass FP4’s dynamic range and enable accurate simulation.
@@
We adopt the widely recognized LLaMA 2 model (Touvron et al., 2023) as the primary model architecture. The training is conducted from scratch using the DCLM dataset (Li et al., 2024a), a comprehensive dataset well-suited for language model pretraining. Hyperparameters remain consistent across precision settings for fair comparison. The learning rate follows a warm-up and cosine decay schedule, with the warm-up phase spanning 5% of total steps and the learning rate gradually decreasing to 10% of its peak over the remaining 90%. The peak learning rate is $3\times 10^{-4}$, with a weight decay of 0.1. For the Adam optimizer, we use $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. For special hyperparameters used in FP4 method, we use $k=5$ for differentiable gradient estimator and select $\alpha=0.99$ as the activation clamp and compensation quantile. Input sequences are fixed at 2048 tokens, and the batch size is 2048, comprising approximately 4M tokens.
@@ 4.2 Main Results
In addition to training loss, we evaluate the models on a diverse set of downstream tasks datasets in a zero-shot manner, including Arc (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), LogiQA (Liu et al., 2020), PiQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), OpenbookQA (ObQA) (Mihaylov et al., 2018), and Lambada (Paperno et al., 2016). These results are obtained through the widely used lm-evaluation-harness library [^1] https://github.com/EleutherAI/lm-evaluation-harness (Gao et al., 2024). As presented in Table 2, models pre-trained with FP4 demonstrate competitive performance in intrinsic in-context learning capabilities. Under the same model size, the average accuracy of FP4-trained models is comparable to, or even slightly exceeds, that of BF16-trained models. Additionally, the results follow the general trend: larger models achieve higher accuracy under the same number of training tokens.
@@ 4.3 Ablation Study
Activation. For activation-only 4-bit quantization (W8A4), we evaluate our Outlier Clamp and Compensation (OCC) method alone against direct quantization. Figure 6(c) reveals that directly quantizing activations in FP4 results in curve divergence, where the loss values turn into NaN (Not a Number) after certain training steps. Outlier clamping and compensation effectively reduces this loss gap, ensuring a good convergence. This experiment re-emphasizes the importance of appropriate treatment of outliers in the absmax quantization framework. For the hyperparameter $\alpha$ in this method, a larger $\alpha$ implies a stronger compensation, but at an increased computational cost. Figure 6(c) shows the model loss under three settings $\alpha=0.999,0.99,0.97$, corresponding to the non-zero elements of the sparse compensation matrix of 0.2%,2% and 6%, respectively. Although experiments show that a higher $\alpha$ leads to better model accuracy, which is consistent with the conclusion of Table 1, we believe that $\alpha=0.99$ is a better choice for comprehensive computational performance considerations.
@@ 5 Related Work
Quantized Training and Inference.When discussing the quantization of large language models (LLMs) for training, we typically refer to Fully Quantized Training (FQT). Related research efforts have generally used Mixed Precision Training (Micikevicius et al., 2017; Mellempudi et al., 2019) frameworks to accelerate model training while maintaining model accuracy. While previous research has mainly concentrated on CNNs or DNNs(Sun et al., 2019; Wang et al., 2018; Banner et al., 2018; Yang et al., 2020), recent studies have demonstrated the feasibility of low-bit mixed precision training for LLMs (Peng et al., 2023; Nvidia, 2022; Fishman et al., 2024; Xi et al., 2024). In contrast to the FQT scheme, research on low-bit computation for inference has focused on Post-Training Quantization (PTQ) and Quantization Aware Training (QAT). While PTQ directly quantizes pre-trained models for inference (Dettmers et al., 2022; Frantar et al., 2022; Lin et al., 2024a; Xiao et al., 2023; Yao et al., 2022; Liu et al., 2024), QAT involves fine-tuning or pre-training the model for better low-bit inference performance (Liu et al., 2023b; Cheng et al., 2023; Wang et al., 2023; Dettmers et al., 2024). Our method differs from QAT, as we aim to accelerate the training process while maintaining performance, rather than solely focusing on improving inference efficiency without consideration for the training speed.
4-bit Quantization. Recent works in PTQ and QAT have successfully applied 4-bit, 2-bit or even 1-bit quantization to LLM inference (Dettmers & Zettlemoyer, 2023; Wu et al., 2023). However, these methods focused on LLM inference, requiring additional computation like calibration set fine-tuning (Wang et al., 2024), rotary matrix and low-rank compensation (Lin et al., 2024b; Ashkboos et al., 2024; Li et al., 2024b), quantization parameters searching (Liu et al., 2023a), or even retraining the whole network (Ma et al., 2024). In the field of FQT, an early study (Sun et al., 2020) applied a 4-bit radix-4 FP4 format to convolutional neural networks (CNNs). MXFP (Rouhani et al., 2023b) introduced a novel quantization data for GPT-style models, but lacked feasibility validation on full FP4 settings. (Xi et al., 2023) proposed an INT4 training framework, but their focus was on fine-tuning tasks with limited applicability to LLM pretraining. In contrast, our work is the first to propose an FP4 training framework tailored for LLMs, validated from scratch, and designed to align with next-generation hardware like Nvidia’s B-series GPUs.
@@
Handling Outliers. Our method for handling activation outliers in LLMs differs significantly from existing approaches, which mainly target model inference (Liu et al., 2023a; Li et al., 2024b; Ashkboos et al., 2024; Liu et al., 2024). Activation outliers in LLMs are typically channel-specific (Xiao et al., 2023; Wei et al., 2022). Channel-wise quantization would reduce quantization loss but conflicts with the computation structure of matrix multiplication in linear layers (Xi et al., 2024; Lee et al., 2024). Previous strategies to solve this problem like smoothing outliers (Xiao et al., 2023) or using rotary matrices (Ashkboos et al., 2024; Liu et al., 2024) rely on offline pre-processing, making them incompatible with pretraining tasks. In contrast, our method addresses outliers dynamically during real-time training without requiring separate calibration datasets, which is critical for maintaining efficiency in pretraining large models.
@@ 7 Conclusion
This work demonstrates the feasibility of using ultra-low precision formats like FP4 for training large language models, offering a pathway toward energy conservation and reduced carbon emissions in AI development. By significantly lowering computational and memory demands, FP4-based methods can democratize access to advanced AI systems while promoting environmental sustainability.
@@ Appendix A Implementation of FP4 quantizaiton
Appendix B Supplementary Proof for Differentiable Quantization Estimator

### v2

@@ 1 Introduction
We conduct comprehensive experiments to demonstrate that our FP4 training framework achieves accuracy comparable to models trained in BF16 or FP8 formats with the same hyperparameters. Leveraging the FP8 tensor cores of NVIDIA H100 GPUs to emulate FP4 computations, we train LLMs with up to 13B parameters and 100B training tokens, with minor training loss gap. For zero-shot evaluation on downstream tasks, model trained with FP4 show competitive results against BF16 models. We anticipate better speed performance gains with the availability of next-generation hardware like NVIDIA’s B-series GPUs. Our training framework can be found at aka.ms/MS.AMP.
@@ 3.1 Differentiable Gradient Estimator
Suppose we quantize the model weight $W$ with a non-differentiable quantization function $f$: $W_{q}=f(W)$. Considering the backward gradient computation for a linear function with quantized weight, the forward pass can be expressed as:
@@
$f(x)=\frac{\delta}{2}\cdot\big{(}1+\text{sign}(\frac{2x}{\delta}-1)\cdot|\frac{2x}{\delta}-1|^{\frac{1}{k}}\big{)}$ (7)
Figure 3(a) illustrates this function under $k=5$ for the range $[0,0.5]$, which represents the first positive quantization interval in the E2M1 quantization scheme. This figure also shows that under the assumption of STE, forward quantization function is equivalent to $f(x)=x$ because $f^{\prime}(x)\equiv 1$. In Equation 7, $\delta$ represents the quantization interval, and $k$ is a parameter that controls the degree of approximation. As $k$ increases, the function curve becomes sharper and more closely resembles the behavior of the original hard quantization function. For details on how Equation 7 is specified and derived, please refer to Appendix C.
@@
$f^{\prime}(x)=\frac{1}{k}\cdot|\frac{2x}{\delta}-1|^{\frac{1}{k}-1}$ (8)
Figure 3(b) and Figure 3(c) show the complete quantization curve $f(x)$ and its derivative $f^{\prime}(x)$ under $k=5$ within the full E2M1 quantization framework. This framework consists of 14 distinct quantization intervals. In practice, the magnitude of $f^{\prime}(x)$ is capped at 3.0 to prevent infinite gradient spikes at $\delta/2$ point, impacting only a very small subset of elements. For the mathematical soundness of this operation, as well as the supplementary integration process and proof for the DGE method in actual training process, please refer to Appendix C.
In practical model training, the Differentiable Gradient Estimator (DGE) is seamlessly integrated into the process. During the forward pass, we retain the hard quantization function for computational efficiency. For the backward pass, a correction term derived from Equation 8 is applied to the weight gradient calculation following Equation 6.
@@ 3.2 Outlier Clamping and Compensation
To address this issue, we propose the Outlier Clamping and Compensation method (OCC) to restrict the range of activation tensors and mitigate the underflow problem. Specifically, we identify outliers—values with the largest absolute magnitudes—through quantile identification and clamp them to a predefined threshold. Given a pre-defined quantile $\alpha$, the clamping function can be expressed as:
@@ 4.1 Experiment Setup
During LLM training, General Matrix Multiplication (GeMM) accounts for over 95% of the computational workload, with this proportion increasing for larger models. Consistent with prior works (Xi et al., 2023; Yang et al., 2020; Dettmers et al., 2022), we focus on 4-bit quantization for GeMM operations, a core feature of FP4 tensor cores. In a GeMM computation $Y=AW$, where $A$ (sequence length × input channels) is the activation tensor and $W$ (input channels × output channels) is the weight tensor, quantization is applied along distinct dimensions to align with matrix multiplication logic: $A$ is quantized token-wise (sequence length dimension), while $W$ is quantized channel-wise (output channels dimension). The aforementioned accuracy-preserving techniques are integrated to minimize quantization error. Since FP4 Tensor Cores are unavailable, we validate FP4 performance using Nvidia H-series GPUs’ FP8 Tensor Cores, which encompass FP4’s dynamic range and enable accurate simulation.
@@
We adopt the widely recognized LLaMA 2 model (Touvron et al., 2023) as the primary model architecture. The training is conducted from scratch using the DCLM dataset (Li et al., 2024), a comprehensive dataset well-suited for language model pretraining. Hyperparameters remain consistent across precision settings for fair comparison. The learning rate follows a warm-up and cosine decay schedule, with the warm-up phase spanning 5% of total steps and the learning rate gradually decreasing to 10% of its peak over the remaining 90%. The peak learning rate is $3\times 10^{-4}$, with a weight decay of 0.1. For the Adam optimizer, we use $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. For special hyperparameters used in FP4 method, we use $k=5$ for differentiable gradient estimator and select $\alpha=0.99$ as the activation clamp and compensation quantile. Input sequences are fixed at 2048 tokens, and the batch size is 2048, comprising approximately 4M tokens.
@@ 4.2 Main Results
In addition to training loss, we evaluate the models on a diverse set of downstream tasks datasets in a zero-shot manner, including Arc (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), LogiQA (Liu et al., 2021), PiQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), OpenbookQA (ObQA) (Mihaylov et al., 2018), and Lambada (Paperno et al., 2016). These results are obtained through the widely used lm-evaluation-harness library [^1] https://github.com/EleutherAI/lm-evaluation-harness (Gao et al., 2024). As presented in Table 2, models pre-trained with FP4 demonstrate competitive performance in intrinsic in-context learning capabilities. Under the same model size, the average accuracy of FP4-trained models is comparable to, or even slightly exceeds, that of BF16-trained models. Additionally, the results follow the general trend: larger models achieve higher accuracy under the same number of training tokens.
Table 3: Perplexity evaluation for downstream tasks between BF16 models and FP4 models under different model sizes.
Size Precision Average Lbd.OAI Lbd.std Pile10k Wikitext
1.3B BF16 37.38 14.98 25.10 82.77 26.65
FP4(Ours) 36.86 15.33 23.07 82.52 26.51
7B BF16 35.06 14.34 23.33 77.72 24.86
FP4(Ours) 35.62 14.29 24.42 78.42 25.36
13B BF16 33.69 12.42 22.45 75.06 24.81
FP4(Ours) 33.99 13.67 21.62 75.84 24.83
Table 3 further presents the perplexity (PPL) evaluation results for several downstream datasets including Lambada OpenAI (Lbd.OAI), Lambada standard (Lbd.std) (Paperno et al., 2016), the Pile 10k (Gao et al., 2020) and Wikitext (Merity et al., 2017). The results demonstrate that FP4 models achieve comparable or even slightly lower PPL than BF16 models. As expected, larger models achieve lower perplexity under the same training token budget.
@@ 4.3 Ablation Study
Activation. For activation-only 4-bit quantization (W8A4), we evaluate our Outlier Clamp and Compensation (OCC) method alone against direct quantization. Figure 6(c) reveals that directly quantizing activations in FP4 results in curve divergence, where the loss values turn into NaN (Not a Number) after certain training steps. Outlier clamping and compensation effectively reduces this loss gap, ensuring a good convergence. This experiment re-emphasizes the importance of appropriate treatment of outliers in the absmax quantization framework. For the hyperparameter $\alpha$ in this method, a smaller $\alpha$ implies a stronger compensation, but at an increased computational cost. Figure 6(c) shows the model loss under three settings $\alpha=0.999,0.99,0.97$, corresponding to the non-zero elements of the sparse compensation matrix of 0.2%,2% and 6%, respectively. Although experiments show that a smaller $\alpha$ leads to better model accuracy, which is consistent with the conclusion of Table 1, we believe that $\alpha=0.99$ is a better choice for comprehensive computational performance considerations.
@@ 5 Related Work
Quantized Training and Inference.When discussing the quantization of large language models (LLMs) for training, we typically refer to Fully Quantized Training (FQT). Related research efforts have generally used Mixed Precision Training (Micikevicius et al., 2017; Mellempudi et al., 2019) frameworks to accelerate model training while maintaining model accuracy. While previous research has mainly concentrated on CNNs or DNNs(Sun et al., 2019; Wang et al., 2018; Banner et al., 2018; Yang et al., 2020), recent studies have demonstrated the feasibility of low-bit mixed precision training for LLMs (Peng et al., 2023; Nvidia, 2022; Fishman et al., 2025; Xi et al., 2024). In contrast to the FQT scheme, research on low-bit computation for inference has focused on Post-Training Quantization (PTQ) and Quantization Aware Training (QAT). While PTQ directly quantizes pre-trained models for inference (Dettmers et al., 2022; Frantar et al., 2023; Lin et al., 2024a; Xiao et al., 2023; Yao et al., 2022; Liu et al., 2024), QAT involves fine-tuning or pre-training the model for better low-bit inference performance (Liu et al., 2023b; Cheng et al., 2023; Wang et al., 2023; Dettmers et al., 2023). Our method differs from QAT, as we aim to accelerate the training process while maintaining performance, rather than solely focusing on improving inference efficiency without consideration for the training speed.
4-bit Quantization. Recent works in PTQ and QAT have successfully applied 4-bit, 2-bit or even 1-bit quantization to LLM inference (Dettmers & Zettlemoyer, 2023; Wu et al., 2023). However, these methods focused on LLM inference, requiring additional computation like calibration set fine-tuning (Wang et al., 2024), rotary matrix and low-rank compensation (Lin et al., 2024b; Ashkboos et al., 2024; Li et al., 2025), quantization parameters searching (Liu et al., 2023a), or even retraining the whole network (Ma et al., 2024). In the field of FQT, an early study (Sun et al., 2020) applied a 4-bit radix-4 FP4 format to convolutional neural networks (CNNs). MXFP (Rouhani et al., 2023b) introduced a novel quantization data for GPT-style models, but lacked feasibility validation on full FP4 settings. (Xi et al., 2023) proposed an INT4 training framework, but their focus was on fine-tuning tasks with limited applicability to LLM pretraining. In contrast, our work is the first to propose an FP4 training framework tailored for LLMs, validated from scratch, and designed to align with next-generation hardware like Nvidia’s B-series GPUs.
@@
Handling Outliers. Our method for handling activation outliers in LLMs differs significantly from existing approaches, which mainly target model inference (Liu et al., 2023a; Li et al., 2025; Ashkboos et al., 2024; Liu et al., 2024). Activation outliers in LLMs are typically channel-specific (Xiao et al., 2023; Wei et al., 2022). Channel-wise quantization would reduce quantization loss but conflicts with the computation structure of matrix multiplication in linear layers (Xi et al., 2024; Lee et al., 2024). Previous strategies to solve this problem like smoothing outliers (Xiao et al., 2023) or using rotary matrices (Ashkboos et al., 2024; Liu et al., 2024) rely on offline pre-processing, making them incompatible with pretraining tasks. In contrast, our method addresses outliers dynamically during real-time training without requiring separate calibration datasets, which is critical for maintaining efficiency in pretraining large models.
@@ 7 Conclusion
This work demonstrates the feasibility of using ultra-low precision formats like FP4 for training large language models, offering a pathway toward energy conservation and reduced carbon emissions in AI development. By significantly lowering computational demand, FP4-based methods can democratize access to advanced AI systems while promoting environmental sustainability.
@@ Appendix A Implementation of FP4 Quantizaiton
Appendix B Theoretical Analysis on Speed Performance and Overhead
To assess the theoretical performance benefit of FP4 acceleration in our mixed-precision training framework, we analyze the theoretical speedup at the level of individual Transformer layer components. Specifically, we compute the floating-point operations (FLOPs) required for each subcomponent in both FP32 and FP4 precision. Let the hidden size be $h$, batch size $b$, and sequence length $s$. The FLOP breakdown per Transformer layer is summarized in Table 5.
Table 5: FLOP breakdown of a standard Transformer layer under FP4 acceleration.
Component Subcomponent FLOPs (FP32) FLOPs (FP4) Speedup Factor
Input LayerNorm $4bsh$ $4bsh$
Multi-Head Attention Query, Key, Value Projections $6bsh^{2}$ $1.5bsh^{2}$
Attention Scores Computation $4bs^{2}h$ $4bs^{2}h$
Softmax Computation $bs^{2}h$ $bs^{2}h$
Output Projection $2bsh^{2}$ $0.5bsh^{2}$
Post-Attention LayerNorm $4bsh$ $4bsh$
Feed-Forward Network (FFN) Up Projection $8bsh^{2}$ $2bsh^{2}$
GeLU Activation $28bsh$ $28bsh$
Down Projection $8bsh^{2}$ $2bsh^{2}$
Total $24bsh^{2}+5bs^{2}h+36bsh$ $6bsh^{2}+5bs^{2}h+36bsh$
Given that the backward pass typically incurs approximately twice the FLOPs of the forward pass, the total computational cost for both forward and backward passes is three times that of the forward pass alone. Thus, the theoretical speedup factor from FP4 acceleration (excluding any overhead) can be expressed as:
$\frac{3\times(24bsh^{2}+5bs^{2}h+36bsh)}{3\times(6bsh^{2}+5bs^{2}h+36bsh)}=\frac{24h+5s+36}{6h+5s+36}$
For a representative case of a 7B model with hidden size $h=4096$ and sequence length $s=2048$, the resulting theoretical speedup is approximately $\dfrac{24\times 4096+5\times 2048+36}{6\times 4096+5\times 2048+36}=3.12$.
In addition to the theoretical speedup from FP4 compute, we also account for two primary sources of computational overhead introduced by our proposed methodologies: Differentiable Gradient Estimator (DGE) and Outlier Clamp and Compensation (OCC).
DGE Overhead. DGE introduces an additional nonlinear function in the backward pass of GEMM operations used for weight updates. As shown in Equation 8, this adds approximately 8 FLOPs per input element. Accumulating over all relevant GEMM operations—including attention query/key/value projections, attention output projection, and MLP up and down projections—the total FLOP overhead per iteration is:
$8\times(3bsh+bsh+4bsh+4bsh)=96bsh$
This overhead occurs only once per forward-backward iteration.
OCC Overhead. OCC introduces extra FP8 sparse matrix multiplications during outlier compensation. Given an activation outlier matrix sparsity of $2(1-\alpha)$, these sparse GEMMs are applied to each of the four primary GEMM computations in a Transformer block, resulting in an additional $2(1-\alpha)\times(12bsh^{2})$ FLOPs per iteration. In our setup, we choose $\alpha=0.99$ to maintain high sparsity in the $\Delta Y$ matrix. While the computational FLOP overhead remains small due to high sparsity, hardware inefficiencies in sparse GEMMs make a high value of $\alpha$ essential for runtime performance.
Adjusted Speedup Estimate. Taking both DGE and OCC overheads into account, the revised theoretical speedup becomes:
$\frac{3\times(24bsh^{2}+5bs^{2}h+36bsh)}{3\times(6bsh^{2}+5bs^{2}h+36bsh+2(1-\alpha)(12bsh^{2}))+96bsh}=\frac{24h+5s+36}{6h+24(1-\alpha)h+5s+68}$
Substituting $h=4096$, $s=2048$, and $\alpha=0.99$, we obtain $\dfrac{24\times 4096+5\times 2048+36}{6\times 4096+24\times 0.01\times 4096+5\times 2048+68}=2.95$
Overhead Impact. The DGE overhead accounts for ${32}/({6h+5s+36})=0.1\%$ and the OCC overhead contributes: ${24(1-\alpha)h}/({6h+5s+36})=5.6\%$ of the total computation. These result in a modest reduction of the ideal FP4 speedup from 3.12 to 2.95, which we consider an acceptable trade-off between computational efficiency and model accuracy.
Appendix C Supplementary Proof for Differentiable Quantization Estimator
C.1 The Derivation of Differentiable Quantization Function
In Section 3.1, we introduce a differentiable function to simulate the quantization funciton in Equation 7. Here we describe in detail how this formula is derived step by step.
To construct a smooth and differentiable approximation to the quantization function, we begin with the power function $f(x)=x^{a}$, where $10<a<$ 1. This function is particularly suitable due to its saturating behavior: it approaches 1 as $x\to 1$, and rapidly decays toward 0 as $x\to 0$. These properties align well with the behavior of the right half of a typical quantization function. To extend the function to the entire real line while maintaining central symmetry around the origin, we first symmetrize it by applying the sign function. This gives:
$f(x)=\text{sign}(x)\cdot|x|^{a}$ (10)
This symmetrized function now resembles a ”soft step” centered at the origin, increasing smoothly from $-1$ to $+1$. We first scale the x-axis adjust the quantization interval from $[-1,1]$ to $[-\delta/2,\delta/2]$, which we need to scale $x$ by $\delta/2$, resulting in:
$f(x)=\text{sign}(\frac{2x}{\delta})\cdot|\frac{2x}{\delta}|^{a}$ (11)
To shift the center of this function from the origin to the midpoint of the quantization interval $[0,\delta]$, we perform a horizontal (x-axis) translation. Specifically, we translate the input by $\delta/2$, which centers the function at $x=\delta/2$. This gives:
$f(x)=\text{sign}\left(\frac{2(x-\frac{\delta}{2})}{\delta}\right)\cdot\left|\frac{2(x-\frac{\delta}{2})}{\delta}\right|^{a}\\ =\text{sign}\left(\frac{2x}{\delta}-1\right)\cdot\left|\frac{2x}{\delta}-1\right|^{a}$ (12)
This modified function now transitions smoothly from negative to positive around the midpoint of the quantization interval, rather than around zero. Similarly, to ensure the output(y) range of the function maps to the desired quantization interval $[0,\delta]$, we apply a vertical (y-axis) translation and scaling. Specifically, we shift the function upward by 1 and then scale it by $\delta/2$, resulting in:
$f(x)=\frac{\delta}{2}\cdot\left(1+\text{sign}\left(\frac{2x}{\delta}-1\right)\cdot\left|\frac{2x}{\delta}-1\right|^{a}\right)$ (13)
This transformation maps the function output to the interval $[0,\delta]$, with a steep but smooth transition occurring around $x=\delta/2$.
Finally, to make the steepness of the approximation more intuitively controllable, we reparameterize the exponent $a$ as the reciprocal of a positive constant $k>1$, i.e., $a=1/k$. Larger values of $k$ lead to steeper transitions, making the function better approximate a hard quantization step, while smaller values of $k$ result in smoother transitions.
C.2 Proof of DGE with Vector-wise Scaling Factors